# SLED: A Speculative LLM Decoding Framework for Efficient Edge Serving

### v4

@@ Abstract
The growing gap between the increasing complexity of large language models (LLMs) and the limited computational budgets of edge devices poses a key challenge for efficient on-device inference, despite gradual improvements in hardware capabilities. Existing strategies, such as aggressive quantization, pruning, or remote inference, trade accuracy for efficiency or lead to substantial cost burdens. This position paper introduces a new framework that leverages speculative decoding, previously viewed primarily as a decoding acceleration technique for autoregressive generation of LLMs, as a promising approach specifically adapted for edge computing by orchestrating computation across heterogeneous devices. We propose SLED, a framework that allows lightweight edge devices to draft multiple candidate tokens locally using diverse draft models, while a single, shared edge server verifies the tokens utilizing a more precise target model. To further increase the efficiency of verification, the edge server batch the diverse verification requests from devices. This approach supports device heterogeneity and reduces server-side memory footprint by sharing the same upstream target model across multiple devices. Our initial experiments with Jetson Orin Nano, Raspberry Pi 4B/5, and an edge server equipped with 4 Nvidia A100 GPUs indicate substantial benefits: ×2.2 more system throughput, ×2.8 more system capacity, and better cost efficiency, all without sacrificing model accuracy.
Index Terms:
@@
Speculative Decoding, Large Language Models, Edge Computing, SLED, Distributed Inference, Token Verification, Model Partitioning, Resource-Aware Serving
I Introduction
LLMs have revolutionized various domains, demonstrating remarkable capabilities in natural language understanding, generation, and complex reasoning [1, 2, 3, 4, 5]. Their widespread adoption has led to transformative applications in areas such as intelligent chatbots, content creation, code generation, and scientific discovery. However, the immense memory and compute footprint associated with state-of-the-art LLMs, often comprising billions or even trillions of parameters, pose significant challenges for deployment. These models typically demand powerful accelerators like GPUs and substantial memory, limiting their direct execution on resource-constrained devices.
@@ I Introduction
Deploying LLMs at the edge, closer to data sources and end-users, offers significant advantages including reduced latency, enhanced privacy, and lower bandwidth consumption[6]. Nevertheless, edge environments, characterized by limited memory, processing power, and energy budgets, present formidable obstacles to efficient LLM inference. Existing strategies to address these limitations include aggressive model compression techniques such as quantization  [7, 8], pruning [9, 10], and knowledge distillation  [11]. Other approaches involve distributed inference, where model layers are partitioned across multiple devices or between edge and cloud[12, 13], or full remote inference, where the entire computation is offloaded to a powerful central server [14, 15]. While these methods show some potential, they often come with trade-offs: compression can sacrifice model accuracy, distributed inference introduces synchronization overheads and is incompatible with heterogeneous edge devices, and remote inference negates the benefits of edge deployment, incurring non-negligible costs.
@@
Figure 1: System overview of the proposed speculative LLM decoding framework for efficient edge serving.
@@
Speculative decoding [16] is a decoding acceleration technique that first generates multiple draft tokens using a relatively small draft model and then verifies them in a single forward pass using a larger, more accurate target model. By generating multiple tokens with the smaller model and validating them in a single pass on the larger model, speculative decoding significantly reduces the number of forward passes required on the large model, thereby accelerating the decoding process. This position paper introduces SLED, shown in Fig. 1, a novel approach that re-imagines speculative decoding as a paradigm specifically tailored for efficient LLM inference at the edge, by intelligently orchestrating computation across heterogeneous devices, as shown in Fig.2. Within the defined service area consisting of an edge server and multiple heterogeneous edge devices, and each edge device is equipped with its own lightweight LLMs scaled according to individual computational and memory resource capacities. These edge devices are responsible for serving diverse LLM-based applications such as intelligent personal assistants, text generation, and semantic analysis, among other tasks. Concurrently, a single, shared edge server, equipped with a more precise target model, efficiently batches and verifies these drafted tokens. The advantages of the SLED are three-fold:
  1. 1.
@@
  2. 2.
    Compared with inference solely on the edge server, SLED reduces the monetary cost for edge users by limiting their usage of server resources–requiring only token verification rather than full generation.
  3. 3.
@@
We compare the system capacity, the number of edge devices supported by the system, of SLED and a centralized serving system with the same response rate but different device types. From Tab. I, we observe that compared with a centralized LLM serving system for the edge, the proposed SLED is capable of increasing the system capacity by 2.6 to 2.9 times.
TABLE I: Capacity of SLED and centralized serving system
@@
  • We propose SLED, a novel speculative decoding framework specifically designed for heterogeneous edge computing environments, enabling efficient LLM inference without accuracy degradation.
@@
The remainder of this paper is organized as follows: Section II reviews existing work in LLM inference for edge computing. Section III details the architectural design and key components of SLED. Section IV presents our experimental setup and discusses the evaluation results. Finally, Section V concludes the paper and outlines future research directions.
II Related Work
TABLE II: Comparison of related work; Edge-Serving: Does the system support edge computing?; Heterogenity: Is the heterogenity of edge devices considered in the system design?; Lossless: Whether does the system deliver LLM service without any performance degradation? Scalable Model: Is the system capable of scaling model according to conditions without too much overhead?
System Edge-Serving Heterogenity Lossless Scalable Model
EdgeShard[12]
Galaxy[13]
Orca[17]
vLLM[18]
FastServe[19]
AWQ[20]
MobileBERT[21]
@@ II Related Work
Model Compression and Lightweight Architectures. To enable LLMs to run on resource-constrained devices, significant efforts have been directed towards model compression. Quantization reduces the numerical precision of model parameters and activations to decrease memory footprint and accelerate computation [22, 23, 20]. Pruning identifies and removes redundant connections or neurons from the neural network without significant performance loss, resulting in sparser and smaller models [24]. Knowledge distillation involves training a smaller ”student” model to mimic the behavior of a larger ”teacher” model, thereby transferring knowledge and achieving comparable performance with a significantly smaller footprint [11]. Beyond these optimization techniques, research has also focused on designing inherently lightweight transformer architectures that are more efficient from the ground up, such as MobileBERT [21], Mamba [25] or other compact variants, often by optimizing attention mechanisms or reducing the number of layers and hidden dimensions. Despite their advantages in reducing model size and computational demands, a common limitation of these model compression techniques is the inherent trade-off with model quality: aggressive compression often leads to a measurable decrease in accuracy compared to their full-sized counterparts.
@@
Edge-Cloud/Server Offloading and Distributed Inference. Another line of research focuses on distributing LLM computation across multiple devices or partitioning tasks between edge and cloud/server infrastructure. Model partitioning schemes divide a large LLM into smaller sub-models, with different parts executed on different devices [12, 26, 13]. For example, EdgeShard[12] partitioned LLM into shards and deploy on distributed devices to benefit from the collaboration among edge devices and cloud server. This often involves pipeline parallelism or tensor parallelism techniques, where different stages or segments of the model’s computation are assigned to different devices. While this allows larger models to run on resource-constrained setups, it introduces communication overheads and synchronization challenges, particularly for heterogeneous hardware and varying network conditions. Edge-cloud offloading dynamically decides which parts of the inference task should be performed locally at the edge and which should be offloaded to a more powerful cloud server, often based on real-time resource availability, network bandwidth, and latency requirements [12]. These methods aim to balance the benefits of edge processing with the computational power of the cloud, but often require sophisticated orchestration and robust connectivity.
@@
Pure Remote Inference. Pure remote inference, where the entire LLM resides on centralized data-center GPUs, represents a prevalent deployment paradigm due to its simplicity and centralized resource utilization. Recent research primarily focuses on resource efficiency and latency optimization. Kwon et al.[18] proposed vLLM with a PagedAttention allocator, significantly reducing KV-cache overhead and fragmentation, achieving up to 4× throughput improvement. Wu et al. [19] introduced FastServe, leveraging multi-level feedback queue scheduling and proactive KV-cache management to reduce tail latency by up to 31× at the 99th percentile. Rajbhandari et al. [27] developed DeepSpeed Inference, combining multiple parallelism strategies with NVMe and CPU off-loading, allowing inference of substantially larger models and reducing latency by up to 7.3×. Crucially, the standard decoding process in remote inference is often autoregressive, generating one token at a time, which can be memory-intensive due to large key-value caches and lead to resource under-utilization on powerful servers. Moreover, the cost associated with cloud GPU instances for continuous, often under-utilized, inference also presents a substantial economic burden, especially for high-throughput scenarios.
@@
Table. II compares the SLED and related works that deliver LLM inference service or propose model variants for edge devices, among which SLED stands out as the only approach that enables lossless LLM inference for heterogeneous edge devices, while maintaining a collaborative design that can flexibly accommodate increasingly large models. SLED directly addresses these limitations by fundamentally altering the inference paradigm. Instead of autoregressive token generation on the powerful central server, SLED offloads the preliminary token drafting to lightweight edge devices. This allows the central server to focus its considerable resources primarily on the more efficient and batchable task of verifying multiple drafted tokens. By doing so, SLED significantly improves the utilization of expensive server-side GPU resources without sacrificing model accuracy, leading to a more cost-effective and scalable distributed LLM inference system.
III SLED Design
Figure 2: System overview of the proposed speculative LLM decoding framework for efficient edge serving.
@@ III SLED Design
III-A Dynamic Drafting on Edge Devices
@@
Previous studies [28], validated by our preliminary experimental results, have established a correlation between the acceptance rate of draft tokens and their associated confidence scores derived from the output logits. As illustrated in Fig.3, draft tokens with higher confidence scores exhibit a significantly increased likelihood of acceptance by the target model.
@@
$c_{s}^{i}\begin{cases}<c_{th},request\;verification\\ \geq c_{th},generate\;another\;token\\ \end{cases},$ (1)
@@
Figure 3: System overview of the proposed speculative LLM decoding framework for efficient edge serving.
III-B Batched Verification on Edge Server
@@
A critical advantage of SLED lies in the target model’s ability to accept and verify draft tokens generated by diverse draft LLMs across heterogeneous edge devices. This compatibility effectively mitigates device heterogeneity, enabling each device to select a draft model suited to its computational constraints while ensuring scalable and efficient inference across a wide range of edge hardware.
IV Evaluation
@@
As for the device setting, we tested Raspberry Pi 4b, 5 and NVIDIA Jetson Orin Nano on the SLED system supported by 4 A100 GPUs.
IV-A Whole System Token Generation Rate (WSTGR)
Figure 4: WSTGR comparison between SLED and centralized LLM serving systems, highlighting improved scalability of the SLED framework.
We first evaluate the Whole System Token Generation Rate (WSTGR), which is defined as the total number of tokens generated and verified by the entire inference system per second, and serves as a metric for the system’s overall productive output[29]. Given a certain time period we measure the total number of tokens generated by the SLED and centralized inference system. The verification workload model is derived from a Raspberry Pi 5 device running a LLaMA 3B model. Additionally, we evaluate two different target models (11B and 70B) on both the SLED system and a centralized inference system. As shown in Fig. 4, for both the 11B and 70B models, the WSTGR increases rapidly in the initial stages as batch size grows, due to the amortization of fixed GPU launch and driver overhead, and improved utilization of GPU cores. The proposed SLED system achieves higher overall throughput than the centralized serving system under identical conditions, including the same number of devices and target model. This observed scalability demonstrates that SLED effectively utilizes distributed edge resources to enhance the system’s token generation capacity.
The more than twofold improvement in WSTGR over the centralized serving system stems from the efficient and balanced distribution of computational tasks in the SLED system. Specifically, in SLED, simple token generation tasks can be handled by relatively small models [16], such as those deployed on edge devices, while more complex tasks are offloaded to larger models on the edge server. This architectural separation allows computation to be distributed across edge devices and the edge server in a more resource-aligned and efficient manner. As a result, the computational capacity of the edge server is reserved for challenging verification tasks, rather than being consumed by processing simpler tokens from edge devices, unlike in a centralized LLM serving setup.
IV-B Speculative Length vs. throughput and capacity
Figure 5: The Impact of speculative length on system-level capacity and device-level throughput, showing the speculative length should be considered to balance the tradeoff.

### v5

@@ Abstract.
The growing gap between the increasing complexity of large language models (LLMs) and the limited computational budgets of edge devices poses a key challenge for efficient on-device inference, despite gradual improvements in hardware capabilities. Existing strategies, such as aggressive quantization, pruning, or remote inference, trade accuracy for efficiency or lead to substantial cost burdens. This position paper introduces a new framework that leverages speculative decoding, previously viewed primarily as a decoding acceleration technique for autoregressive generation of LLMs, as a promising approach specifically adapted for edge computing by orchestrating computation across heterogeneous devices. We propose SLED, a framework that allows lightweight edge devices to draft multiple candidate tokens locally using diverse draft models, while a single, shared edge server verifies the tokens utilizing a more precise target model. To further increase the efficiency of verification, the edge server batches the diverse verification requests from devices. This approach supports heterogeneous devices and reduces server-side memory footprint by sharing a single upstream target model across devices. Our initial experiments with Jetson Orin Nano, Raspberry Pi 4B/5, and an edge server equipped with 4 Nvidia A100 GPUs indicate substantial benefits: ×2.2 higher system throughput, ×2.8 higher system capacity, and better cost efficiency, all without sacrificing model accuracy.
@@
Speculative Decoding, Large Language Models, Edge Computing, Distributed Inference, Token Verification, Resource-Aware Serving
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM/IEEE Symposium on Edge Computing; December 03–06, 2025; Washington, D.C., USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Distributed computing methodologiesccs: Computing methodologies Distributed artificial intelligenceccs: Computer systems organization Distributed architecturesccs: Computing methodologies Natural language processingccs: Networks Cloud computingccs: General and reference General conference proceedingsjournalyear: 2025copyright: ccconference: The Tenth ACM/IEEE Symposium on Edge Computing; December 3–6, 2025; Arlington, VA, USAbooktitle: The Tenth ACM/IEEE Symposium on Edge Computing (SEC ’25), December 3–6, 2025, Arlington, VA, USAdoi: 10.1145/3769102.3770608isbn: 979-8-4007-2238-7/25/12
[…]
[4.0]by-nc
1. Introduction
@@
LLMs have revolutionized various domains, demonstrating remarkable capabilities in natural language understanding, generation, and complex reasoning (NEURIPS2020_1457c0d6). Their widespread adoption has led to transformative applications in areas such as intelligent chatbots, content creation, code generation, and scientific discovery. However, the immense memory and compute footprint associated with state-of-the-art LLMs, often comprising billions or even trillions of parameters, pose significant challenges for deployment. These models typically demand powerful accelerators like GPUs and substantial memory, limiting their direct execution on resource-constrained devices. Deploying LLMs at the edge, closer to data sources and end-users, offers significant advantages including reduced latency, enhanced privacy, and lower bandwidth consumption(8763885). Nevertheless, edge environments, characterized by limited memory, processing power, and energy budgets, present formidable obstacles to efficient LLM inference. Existing strategies to address these limitations include aggressive model compression techniques such as quantization (MLSYS2024_42a452cb; zeng2024abqllmarbitrarybitquantizedinference), pruning(ma2023llmprunerstructuralpruninglarge; sun2024simpleeffectivepruningapproach), and knowledge distillation (hinton2015distillingknowledgeneuralnetwork). Other approaches involve distributed inference, where model layers are partitioned across multiple devices or between edge and cloud(10818760; ye2025jupiter), or full remote inference, where the entire computation is offloaded to a powerful central server(gao2025collaborativespeculativeinferenceefficient; yu2024edgellmenablingefficientlarge). While these methods show some potential, they often come with trade-offs: compression can sacrifice model accuracy, distributed inference introduces synchronization overheads and is incompatible with heterogeneous edge devices, and remote inference negates the benefits of edge deployment, incurring non-negligible costs.
@@
Figure 1. System overview of the proposed speculative LLM decoding framework for efficient edge serving.:
@@
Speculative decoding (leviathan2023fast) is a decoding acceleration technique that first generates multiple draft tokens using a relatively small draft model and then verifies them in a single forward pass using a larger, more accurate target model. By generating multiple tokens with the smaller model and validating them in a single pass on the larger model, speculative decoding significantly reduces the number of forward passes required on the large model, thereby accelerating the decoding process. This position paper introduces SLED, shown in Fig. 1, a novel approach that re-imagines speculative decoding as a paradigm specifically tailored for efficient LLM inference at the edge, by intelligently orchestrating computation across heterogeneous devices. Within the defined service area consisting of an edge server and multiple heterogeneous edge devices, and each edge device is equipped with its own lightweight LLMs scaled according to individual computational and memory resource capacities. These edge devices are responsible for serving diverse LLM-based applications such as intelligent personal assistants, text generation, and semantic analysis, among other tasks. Concurrently, a single, shared edge server, equipped with a more precise target model, efficiently batches and verifies these drafted tokens.
The advantages of the SLED are threefold:
  1. (1)
@@
  2. (2)
    Compared with inference solely on the edge server, SLED reduces the monetary cost for edge users by limiting their use of server resources, requiring only token verification rather than full generation.
  3. (3)
@@
We compare the system capacity, the number of edge devices supported by the system, of SLED and a centralized serving system with the same response rate but different device types. From Tab. 1, we observe that compared with a centralized LLM serving system for the edge, the proposed SLED is capable of increasing the system capacity by 2.6 to 2.9 times.
Table 1. Capacity of SLED and centralized serving system (Capacity = maximum number of supported devices at target response rate (devices))
@@
  • We propose SLED, a novel speculative decoding framework specifically designed for heterogeneous edge computing environments by batching the verification requests from varying draft models on the server, enabling efficient LLM inference without accuracy degradation.
@@
The remainder of this paper is organized as follows: Section 2 reviews existing work in LLM inference for edge computing. Section 3 details the architectural design and key components of SLED. Section 4 presents our experimental setup and discusses the evaluation results. Finally, Section 5 concludes the paper and outlines future research directions.
2. Related Work
Table 2. Comparison of related work; Edge-Serving: Does the system support edge computing?; Heterogeneity: Is the heterogenity of edge devices considered in the system design?; Lossless: Whether does the system deliver LLM service without any performance degradation? Scalable Model: Is the system capable of scaling model according to conditions without too much overhead?
System Edge-Serving Heterogeneity Lossless Scalable Model
EdgeShard(10818760)
Galaxy(ye2025jupiter)
Orca(yu2022orca)
vLLM(kwon2023vllm)
FastServe(wu2024fastserve)
AWQ(MLSYS2024_42a452cb)
MobileBERT(sun2020mobile)
@@ 2.1. Model Compression and Lightweight Architectures
To enable LLMs to run on resource-constrained devices, significant efforts have been directed towards model compression. Quantization reduces the numerical precision of model parameters and activations to decrease memory footprint and accelerate computation (dettmers2022llmint88bitmatrixmultiplication; Liu_2023; MLSYS2024_42a452cb). Pruning identifies and removes redundant connections or neurons from the neural network without significant performance loss, resulting in sparser and smaller models (han2016deepcompressioncompressingdeep). Knowledge distillation involves training a smaller ”student” model to mimic the behavior of a larger ”teacher” model, thereby transferring knowledge and achieving comparable performance with a significantly smaller footprint (hinton2015distillingknowledgeneuralnetwork). Beyond these optimization techniques, research has also focused on designing inherently lightweight transformer architectures that are more efficient from the ground up, such as MobileBERT (sun2020mobile), Mamba (gu2024mambalineartimesequencemodeling) or other compact variants, often by optimizing attention mechanisms or reducing the number of layers and hidden dimensions. Recent efforts, such as EdgeLLM (10812936), have explored speculative decoding for on-device inference by introducing compute-efficient branch navigation and adaptive fallback strategies to reduce resource demands. However, co-locating both draft and target models on edge devices further strains limited resources, and the output quality remains constrained by the size of the deployable target model. Despite their advantages in reducing model size and computational demands, a common limitation of these model compression techniques is the inherent trade-off with model quality: aggressive compression often leads to a measurable decrease in accuracy compared to their full-sized counterparts.
2.2. Edge-Cloud/Server Offloading and Distributed Inference
@@
Another line of research focuses on distributing LLM computation across multiple devices or partitioning tasks between edge and cloud/server infrastructure. Model partitioning schemes divide a large LLM into smaller sub-models, with different parts executed on different devices (10818760; ye2024galaxyresourceefficientcollaborativeedge; ye2025jupiter). For example, EdgeShard(10818760) partitioned LLM into shards and deploy on distributed devices to benefit from the collaboration among edge devices and cloud server. This often involves pipeline parallelism or tensor parallelism techniques, where different stages or segments of the model’s computation are assigned to different devices. While this allows larger models to run on resource-constrained setups, it introduces communication overheads and synchronization challenges, particularly for heterogeneous hardware and varying network conditions. Edge-cloud offloading dynamically decides which parts of the inference task should be performed locally at the edge and which should be offloaded to a more powerful cloud server, often based on real-time resource availability, network bandwidth, and latency requirements (10818760). These methods aim to balance the benefits of edge processing with the computational power of the cloud, but often require sophisticated orchestration and robust connectivity.
2.3. Pure Remote Inference
@@
Pure remote inference, where the entire LLM resides on centralized data-center GPUs, represents a prevalent deployment paradigm due to its simplicity and centralized resource utilization. Recent research primarily focuses on resource efficiency and latency optimization. Kwon et al.(kwon2023vllm) proposed vLLM with a PagedAttention allocator, significantly reducing KV-cache overhead and fragmentation, achieving up to 4× throughput improvement. Wu et al. (wu2024fastserve) introduced FastServe, leveraging multi-level feedback queue scheduling and proactive KV-cache management to reduce tail latency by up to 31× at the 99th percentile. Rajbhandari et al. (rajbhandari2022deepspeed) developed DeepSpeed Inference, combining multiple parallelism strategies with NVMe and CPU off-loading, allowing inference of substantially larger models and reducing latency by up to 7.3×. Crucially, the standard decoding process in remote inference is often autoregressive, generating one token at a time, which can be memory-intensive due to large key-value caches and lead to resource under-utilization on powerful servers. Moreover, the cost associated with cloud GPU instances for continuous, often under-utilized, inference also presents a substantial economic burden, especially for high-throughput scenarios.
@@
Table 2 compares the SLED and related works that deliver LLM inference service or propose model variants for edge devices, among which SLED stands out as the only approach that enables lossless LLM inference for heterogeneous edge devices, while maintaining a collaborative design that can flexibly accommodate increasingly large models. SLED directly addresses these limitations by fundamentally altering the inference paradigm. Instead of autoregressive token generation on the powerful central server, SLED offloads the preliminary token drafting to lightweight edge devices. This allows the central server to focus its considerable resources primarily on the more efficient and batchable task of verifying multiple drafted tokens. By doing so, SLED significantly improves the utilization of expensive server-side GPU resources without sacrificing model accuracy, leading to a more cost-effective and scalable distributed LLM inference system.
3. SLED Design
Figure 2. SLED architecture and request/verification data flow:
3.1. Speculative Decoding
SLED adopts speculative decoding(leviathan2023fast), a recently proposed paradigm that leverages a lightweight draft model to generate multiple tokens. Instead of waiting for a single token from the large target model at each step, the draft model speculatively predicts a candidate sequence, which is then selectively verified or corrected by the more accurate but slower target model with one single forwarding. This design decouples the fast-path token generation from the slow-path verification, enabling improved throughput and reduced end-to-end latency without compromising output quality. In SLED, this speculative execution is further optimized for edge-server deployment under variable connectivity and resource constraints.
According to the original work on speculative decoding(leviathan2023fast), a draft token $\tilde{x}_{n+1}$ is sampled from the draft model’s output distribution $p(x\mid x_{1},x_{2},\cdots,x_{n})$. This draft token is then accepted with probability:
(1) $\alpha=\min\!\left(1,\frac{p_{\mathrm{target}}(\tilde{x}_{n+1}\mid x_{1},\dots,x_{n})}{p_{\mathrm{draft}}(\tilde{x}_{n+1}\mid x_{1},\dots,x_{n})}\right),$
where $p_{\text{target}}$is the distribution given by the large target model, and $p_{\text{draft}}$ is that of the smaller draft model. Tokens that fail this acceptance criterion are resampled directly from $p_{\text{target}}-p_{\text{{draft}}}$. As a result, the generated sequence strictly follows the target model’s distribution, ensuring no degradation in accuracy compared with standard autoregressive decoding.
3.2. System Overview
@@
3.3. Dynamic Drafting on Edge Devices
@@
Previous studies (huang2025specserve), validated by our preliminary experimental results, have established a correlation between the acceptance rate of draft tokens and their associated confidence scores derived from the output logits. As illustrated in Fig.3, draft tokens with higher confidence scores exhibit a significantly increased likelihood of acceptance by the target model.
@@
(2) $c_{s}^{i}\begin{cases}<c_{th},request\;verification\\ \geq c_{th},generate\;another\;token\\ \end{cases},$
@@
Figure 3. Acceptance rate vs. confidence of draft tokens:
3.4. Batched Verification on Edge Server
@@
A critical advantage of SLED lies in the target model’s ability to accept and verify draft tokens generated by diverse draft LLMs across heterogeneous edge devices provided that all edge devices share the same tokenizer. This compatibility effectively mitigates device heterogeneity, enabling each device to select a draft model suited to its computational constraints while ensuring scalable and efficient inference across a wide range of edge hardware.
4. Evaluation
@@
As for the device setting, we tested Raspberry Pi 4b, 5 and NVIDIA Jetson Orin Nano on the SLED system supported by 4 A100 GPUs. We deployed LLaMA family, including model sizes of 1B, 3B, 11B and 70B on the edge devices and verification server, supporting up to context length of 128k.
4.1. Whole System Token Generation Rate (WSTGR)
Figure 4. WSTGR comparison between SLED and centralized LLM serving systems, highlighting improved scalability of the SLED framework. :
We first evaluate the Whole System Token Generation Rate (WSTGR), which is defined as the total number of tokens generated and verified by the entire inference system per second, and serves as a metric for the system’s overall productive output(liu2024optimizing). Given a certain time period we measure the total number of tokens generated by the SLED and centralized inference system. The verification workload model is derived from a Raspberry Pi 5 device running a LLaMA 3B model. Additionally, we evaluate two different target models (11B and 70B) on both the SLED system and a centralized inference system. As shown in Fig. 4, for both the 11B and 70B models, the WSTGR increases rapidly in the initial stages as batch size grows, due to the amortization of fixed GPU launch and driver overhead, and improved utilization of GPU cores. The proposed SLED system achieves higher overall throughput than the centralized serving system under identical conditions, including the same number of devices and target model. This observed scalability demonstrates that SLED effectively utilizes distributed edge resources to enhance the system’s token generation capacity.
The more than twofold improvement in WSTGR over the centralized serving system stems from the efficient and balanced distribution of computational tasks in the SLED system. Specifically, in SLED, simple token generation tasks can be handled by relatively small models (leviathan2023fast), such as those deployed on edge devices, while more complex tasks are offloaded to larger models on the edge server. This architectural separation allows computation to be distributed across edge devices and the edge server in a more resource-aligned and efficient manner. As a result, the computational capacity of the edge server is reserved for challenging verification tasks, rather than being consumed by processing simpler tokens from edge devices, unlike in a centralized LLM serving setup.
4.2. Speculative Length vs. throughput and capacity
Figure 5. The Impact of speculative length on system-level capacity and device-level throughput, showing the speculative length should be considered to balance the tradeoff.: